OpenStudy (anonymous):

Roberta needs ribbon for a craft project. The ribbon sells for $3.75 per yard. Find the cost, in dollars, for 48 inches of the ribbon.

OpenStudy (anonymous):

first we convert inches to yards 1 yard = 36 inches so 48 inches = 1/36 * 48 = 4/3 yards then we find its cost Cost = 4/3 * 3.75 = $5

OpenStudy (anonymous):

thank you! that was a better explanation then my math teacher!